# The Base-Rate Effect on LLM Benchmark Performance: Disambiguating Test-Taking Strategies from Benchmark Performance

### v1

@@ 1 Introduction
22footnotetext: Equal Contribution
@@ 2 Background & Related Work
The MMLU benchmark aims to jointly measure language understanding and knowledge retrieval abilities. It consists of 15908 multiple choice questions distributed across 57 subject areas. Each question is associated with four answer choices that are unevenly distributed across subjects: A ($\mu$=0.231, $\sigma$=0.042), B ($\mu$=0.245, $\sigma$=0.042), C ($\mu$=0.254, $\sigma$=0.044), D ($\mu$=0.270, $\sigma$=0.078). Models are evaluated on their accuracy at selecting the correct answer choice however the method of selection is not prescribed, with many models reporting 0-shot and/or 5-shot performance as measured by a cloze test prompting.
Despite its popularity, MMLU and similar multiple choice question answering (MCQA) benchmarks have seen criticism. Gema et al. (2024) find numerous factual and catastrophic formatting errors across the MMLU dataset. This serves to lower the expected ideal accuracy but does not diminish the overall ideal roll of the benchmark.
@@
Mizrahi et al. (2023) propose rewording perturbation of prompts in numerous benchmarks, including MMLU, to improve robustness of results. This addresses a deficiency in benchmark results, but not the BRP effect specifically addressed here.
@@ 2.1 Cloze & Counterfactual Prompting
We consider that the BRP may influence the reported metrics and give a skewed perception of model understanding (H2). We evaluate an alternative, semantically equivalent, prompting pattern referred to as counterfactual (CF) prompting. CF prompting moves the target completion into the context and employs a canary completion shared across the new contexts. Model choice is observed, as in the cloze test, by the relative likelihood of the completion ie. $max_{a\in L=\{A,B,C,D\}}P(t|Q,C_{L},a)$, where $t$ is the canary token.
@@ 3 Experimental Design & Results
Here we present the experiments used to evaluate the hypotheses in Table 1 and the associated results. All experiments used an A100 GPU Google Colab environment for ~45 GPU hours. Token likelihoods were obtained using a fork of the minicons Python library Misra (2022).
3.1 General Prompt Design
Cloze prompts all follow the format "QUESTION (A) ANSWER1 (B) ANSWER2 (C) ANSWER3 (D) ANSWER4. Of the answer choices above, the best answer choice is (". Probabilities for all four answer choice labels (A,B,C,D) are measured.
CF prompting moves the labels to the context by changing the final clause to "answer X is the", where X is an answer choice label and "best" is the canary token used to measure answer preference.
@@ 3.2 Base-Rate Probability
MMLU provides few requirements on the prompt format, allowing researchers to adapt the format to the LLM’s unique needs. We define BRP as the likelihood of generating a token given equivalent semantic context in each answer choice with no question text. Practically, BRPs are measured using a set of control prompts that follow the cloze format "Select an answer choice (A) choice (B) choice (C) choice (D) choice. Of the answer choices above, the best answer choice is (". This prompt is also modified to a CF form as described in section 3.1.
Fully empty context is avoided to prevent misinterpreting A as an article, and it allows controlling for ordering effects. The control is generated with all 24 permuted orderings of the answer choice labels. BRP is measured as the average probability of each choice label over all positions.
We investigate how much this biased label BRP affects performance on the MMLU task. We first split the dataset by the correct answer choice and measure accuracy for each resulting subset independently. The results of this are shown in the middle two rows of figure 1, where each dot represents the accuracy on the label-split subset of the data for a single subject. Finally, for each subject, we measure the Pearson’s r correlation between the accuracy for each answer label versus the BRP of that answer label.
@@ 3.2.1 Base-Rate Probability Results
Accuracy measures show a similar strong disparity between answer choices. The most egregious example being GPT-2 under cloze testing which has near perfect accuracy when A is correct and near zero accuracy otherwise, suggesting that the model nearly exclusively answers A regardless of context. This same effect is present, though much less pronounced with Mistral and SOLAR. Correlation measurement shows that these three models correlate to a nearly perfect level across all subjects, suggesting a strong causal link between BRPs and accuracy in cloze tasks. Notably, both accuracy disparity and correlation measured are insignificant in the LLaMa models.
@@ 3.2.2 CF Prompting Does Not Eliminate BRP
3.3 Nvr-X-MMLU Testing
In this section we propose the Nvr-X-MMLU, an MMLU variation designed to disentangle BRP effects and task performance and more accurately report the latter. Nvr-X-MMLU consists of four variations of the MMLU dataset. In each Nvr-X variation, answer choice content is remapped to answer choice labels such that the correct answer content is never assigned to label X. The new correct answer label is chosen uniformly at random from the non-excluded labels and incorrect answer labels are subsequently reassigned arbitrarily.
The performance of the model is measured as the minimum accuracy over the four Nvr-X variation sets. Just as with the MMLU, random guessing on Nvr-X-MMLU results in $25\%$ accuracy, while base-rate driven preference or anti-preference for a specific label will achieve $0\%$ and $22\%$ respectively. This is in contrast to the standard MMLU on which such base-rate driven preference will still achieve $25\%$. The resulting accuracy better measures the model’s understanding and factual knowledge.
@@ 3.3 Nvr-X-MMLU Testing
3.3.1 NVR-X-MMLU Results
@@
When measured via CF test, we find that models that do well on MMLU and Never-X-MMLU with cloze testing often perform poorly on counterfactual variations, with only LLaMa 3 even outperforming random chance. Additionally, all models exhibit significant label preference when measured with CF. Users interact with LLMs using a variety of patterns White et al. (2023). The discrepancy between cloze and CF results suggests that model understanding can be brittle, degrading performance across semantically equivalent tasks based on interaction pattern.
@@ 5 Limitations
6 Ethical Considerations
Strategic behavior is an important part of intelligence but is not what is intended to be measured in the case of the MMLU task. We find the novel Nvr-X-MMLU dataset more accurately measures the intended abilities. That being said, models which are specialized for strategic behavior may perform poorly on the Nvr-X-MMLU dataset. When considering the suitability of a model, users should not take benchmark metrics as definitive measures of generic capability. Instead, they should be understood within context of the task. Though some models performed much more poorly on Nvr-X-MMLU, this does not generically denote that the affected models are of a poor quality.
All software used is open source and was used in accordance with the associated license and the intended use stated or implied above. This includes: minicons Misra (2022), MMLU Hendrycks et al. (2020), GPT2 Radford et al. (2019), LLaMa2 Touvron et al. (2023), LLaMa3 Meta (2024), Mistral Jiang et al. (2023), Solar Kim et al. (2023). The Nvr-X-MMLU test created here is released as open source under MIT license.

### v2

@@ 1 Introduction
22footnotetext: Equal Contribution11footnotetext: https://github.com/KyleAMoore/MMLU-cloze-vs-cf
@@ 2 Background & Related Work
The MMLU benchmark aims to jointly measure language understanding and knowledge retrieval abilities. It consists of 15908 multiple choice questions distributed across 57 subject areas. Each question is associated with four answer choices that are unevenly distributed across subjects: A ($\mu$=0.231, $\sigma$=0.042), B ($\mu$=0.245, $\sigma$=0.042), C ($\mu$=0.254, $\sigma$=0.044), D ($\mu$=0.270, $\sigma$=0.078). Models are evaluated on their accuracy at selecting the correct answer choice. However the method of selection is not prescribed, with many models reporting 0-shot and/or 5-shot performance as measured by a cloze test prompting methods.
Example Shared Context
What element is most common among the Jovian Planets?
(A) Hydrogen (B) Helium (C) Carbon (D) Oxygen.
Of the answer choices above,
Prompting Method
Method-Specific Context(s)
Token(s) Measured
Cloze the best answer choice is ( ____ A
B
C
D
CF
answer choice A is the ____
best
answer choice B is the ____
answer choice C is the ____
answer choice D is the ____
Table 2: Comparison of cloze vs CF prompting methods. Each method measures the probability of each token measured given the shared context and the method-specific contexts. Cloze prompting uses a single method-specific context and measures the probability of multiple candidate tokens. CF prompting uses a different method-specific context for each candidate token and measures the probabilities of the same canary token.
Despite its popularity, MMLU and similar multiple choice question answering (MCQA) benchmarks have seen criticism. Gema et al. (2024) find numerous factual and formatting errors across the MMLU dataset. This may lower the expected accuracy but does not diminish MMLU’s overall role.
@@
Mizrahi et al. (2023) propose rewording perturbation of prompts in numerous benchmarks, including MMLU, to improve robustness of results. This addresses a deficiency in benchmark results, but not the BRP effect specifically addressed here. Concurrent to our work, Wei et al. (2024) proposes techniques for improving performance on tasks that are susceptible to similar base rate effects. Our proposed method differs in that it addresses the effect at task level rather than strategy level.
@@ 2.1 Cloze & Counterfactual Prompting
We consider that the BRP may influence the reported metrics and give a skewed perception of model understanding (H2). We evaluate an alternative, semantically equivalent, prompting pattern referred to as counterfactual (CF) prompting. CF prompting moves the target completion into the context and employs a canary completion shared across the new contexts. Model choice is observed, as in the cloze test, by the relative likelihood of the completion ie. $max_{a\in L=\{A,B,C,D\}}P(t|Q,C_{L},a)$, where $t$ is the canary token. Both prompting methods are exemplified in Table 2.
@@ 3 Experimental Design & Results
Here we present the experiments used to evaluate the hypotheses in Table 1 and the associated results. All experiments used an A100 GPU Google Colab environment for ~45 GPU hours. Token likelihoods were obtained using a fork of the minicons Python library Misra (2022). Prompts used in all experiments use the format exemplified in Table 2.
@@ 3.1 Base-Rate Probability
MMLU provides few requirements on the prompt format, allowing researchers to adapt the format to the LLM’s unique needs. We define BRP as the likelihood of generating a token given equivalent semantic context in each answer choice with no question text. Practically, BRPs are measured using a set of control prompts that follow the cloze format "Select an answer choice (A) choice (B) choice (C) choice (D) choice. Of the answer choices above, the best answer choice is (". This prompt is also modified to a CF form as exemplified in Table 2.
Fully empty context is avoided to prevent misinterpreting the token A as an article. The control is generated with all 24 permuted orderings of the answer choice labels. BRP is measured as the average probability of each choice label over all positions.
We investigate how much this biased label BRP affects performance on the MMLU task. We first split the dataset by the correct answer choice and measure accuracy for each resulting subset independently. The results of this are shown in the middle two rows of Figure 1, where each dot represents the accuracy on the label-split subset of the data for a single subject. Finally, for each subject, we measure the Pearson’s r correlation between the accuracy versus the BRP for each answer label.
@@ 3.1.1 Base-Rate Probability Results
Accuracy measures show a similar strong disparity between answer choices. The most egregious example being GPT-2 Radford et al. (2019) under cloze testing, which has near perfect accuracy when A is correct and near zero accuracy otherwise, suggesting that the model nearly exclusively answers A regardless of context. This same effect is present, though much less pronounced with Mistral Jiang et al. (2023) and SOLAR Kim et al. (2023). Results show that these three models correlate nearly perfectly across all subjects, suggesting a strong causal link between BRPs and accuracy in cloze tasks. Notably, both accuracy disparity and correlation are insignificant in the LLaMa Touvron et al. (2023); Meta (2024) models.
@@ 3.1.2 CF Prompting Does Not Eliminate BRP
3.2 Nvr-X-MMLU
In this section we propose the Nvr-X-MMLU, an MMLU variation designed to disentangle BRP effects and task performance and more accurately report the latter. Nvr-X-MMLU consists of four variations of the MMLU dataset. In each Nvr-X variation, answer choice content is remapped to answer choice labels such that the correct answer content is never assigned to label X. The new correct answer label is chosen uniformly at random from the non-excluded labels and incorrect answer labels are subsequently reassigned arbitrarily. The process is described in Algorithm 1 for the Nvr-A dataset. Nvr-B, Nvr-C, and Nvr-D are defined similarly by changing only the value of $X$.
The performance of the model is measured as the minimum accuracy over the four Nvr-X variation sets. Just as with the MMLU, random guessing on Nvr-X-MMLU results in $25\%$ accuracy, while base-rate driven exclusive preference or complete anti-preference for a specific label will achieve $0\%$ and $0(\frac{1}{3})+\frac{1}{3}(0)+\frac{1}{3}(\frac{1}{3})+\frac{1}{3}(\frac{1}{3})\approx 22\%$, respectively. The latter is due to the probability of an answer choice being correct, in parentheses, and the probability that it is selected, outside of parentheses. This is in contrast to the standard MMLU on which such base-rate driven preference will still achieve $25\%$. The resulting accuracy better measures the model’s understanding and factual knowledge.
@@ 3.2 Nvr-X-MMLU
Algorithm 1 MMLU $\rightarrow$ Nvr-A-MMLU
$Q\leftarrow$ MMLU Questions
$Q^{\overline{X}}\leftarrow[\ ]$triangleright Nvr-A Questions
$X\leftarrow 0$triangleright A=0, B=1, etc.
for all $q\in Q$ do
     $A\leftarrow$ Answer choices for $q_{n}$
     $c\leftarrow$ correct choice index $\in A$
     for $a_{i}\in A$ do
         $A_{i}\leftarrow(a_{i},false)$
     end for
     $A_{i=c}\leftarrow(a_{i=c},true)$
     repeat
         $A\leftarrow shuffle(A)$
     until $A_{i=X}[1]\neq true$
     for $a_{i}\in A$ do
         $A_{i}\leftarrow a[0]$
     end for
     insert $(q,A)$ into $Q^{\overline{X}}$
end for
3.3 NVR-X-MMLU Results
@@
When measured via CF test, we find that models that do well on MMLU and Never-X-MMLU with cloze testing often perform poorly on CF variations, with only LLaMa 3 even outperforming random chance. Additionally, all models exhibit significant label preference when measured with CF. Users interact with LLMs using a variety of patterns White et al. (2023). The discrepancy between cloze and CF results suggests that model understanding can be brittle, degrading performance across semantically equivalent tasks based on interaction pattern.
@@ 4 Discusssion
It is important to note that the models tested (listed below) have an impact on the obtained results. It may be that other models or methodological variations may show BRP effects to greater or lesser extents than are observed here. All software used is open source and was used in accordance with the associated license and the intended use stated or implied above. This includes: minicons Misra (2022), MMLU Hendrycks et al. (2020), GPT2 Radford et al. (2019), LLaMa2 Touvron et al. (2023), LLaMa3 Meta (2024), Mistral Jiang et al. (2023), Solar Kim et al. (2023). The Nvr-X-MMLU test created here is released as open source under MIT license at https://github.com/KyleAMoore/MMLU-cloze-vs-cf.
Ethical Considerations
Strategic behavior like defaulting in the face of uncertainty is an important part of intelligence but is not what is intended to be measured in the case of the MMLU task. We find the novel Nvr-X-MMLU dataset more accurately measures the intended abilities. That being said, models which are specialized for strategic behavior may perform poorly on the Nvr-X-MMLU dataset. When considering the suitability of a model, users should not take benchmark metrics as definitive measures of generic capability. Instead, they should be understood within context of the task. Though some models performed much more poorly on Nvr-X-MMLU, this does not generically denote that the affected models are of a poor quality.